Item1 10points ItemSkipped eBook Print References Check my workCheck My Work button is now enabled1 Item...

Item1

10points

ItemSkipped

eBook

Print

References

Check my workCheck My Work button is now enabled1

Item 1

Item 1 10 points Item Skipped

Wave Runners is a boat rental business that has the following costs over the relevant range of 17,000 to 23,000 operating hours for its boats:

Operating Hours
17,000 19,000 21,000 23,000
Total costs:
Variable costs $59,500
Fixed costs 3,120,180
Total costs $3,179,680 $0 $0 $0
Cost per unit:
Variable cost
Fixed cost
Total cost per hour $0.00 $0.00 $0.00 $0.00

Required:

1. Complete the schedule of the company’s total and unit costs. (Round "Cost per unit" answers to 2 decimal places.)

2. Assume that Wave Runners rents it boats for $168 per hour and had 20,000 hours of rentals last year. Prepare a contribution format income statement for the year.

.
Answer #1
Operating Hours
17,000 19,000 21,000 23,000
Total costs:
Variable costs $59,500 19,000 x 3.5 = 66,500 21,000 x 3.5 = 73,500 23,000 x 3.5 =80,500
Fixed costs 3,120,180 3,120,180 3,120,180 3,120,180
Total costs $3,179,680 $3,186,680 $3,193,680 $3,200,680
Cost per unit:
Variable cost 59,500/17,000 = 3.5 66,500/19,000 = 3.5 73,500/21,000 = 3.5 80,500/23,000 = 3.5
Fixed cost 3,120,180/17,000 = 183.54 3,120,180/19,000 = 164.22 3,120,180/21,000 = 148.58 3,120,180/23,000 = 135.66
Total cost per hour $187.04 $167.72 $152.08 $139.16

Variable cost per unit = Total variable cost / Number of units

Total variable cost = Variable cost per unit x Number of units

Fixed cost per unit = Total fixed cost / Number of units

Total fixed cost does not change due to change in output.

2.

Contribution Format Income statement
Sales ( 20,000 x 168) 3,360,000
Variable cost ( 20,000 x 3.5) 70,000
Contribution Margin 3,290,000
Fixed cost -3,120,180
Operating Income $169,820

Kindly comment if you need further assistance. Thanks‼!